A firm will pay a dividend of $2.01 next year. The dividend is expected to grow at a constant rate of 3.75% forever and the required rate of return is 14.47%. What is the value of the stock?


Submit
Answer format: Currency: Round to: 2 decimal places.


unanswered
not_submitted
#6
Derek plans to retire on his 65th birthday. However, he plans to work part-time until he turns 75.00. During these years of part-time work, he will neither make deposits to nor take withdrawals from his retirement account. Exactly one year after the day he turns 75.0 when he fully retires, he will wants to have $2,760,505.00 in his retirement account. He he will make contributions to his retirement account from his 26th birthday to his 65th birthday. To reach his goal, what must the contributions be? Assume a 5.00% interest rate.


Submit
Answer format: Currency: Round to: 2 decimal places.


unanswered
not_submitted
#7
A bank offers 8.00% on savings accounts. What is the effective annual rate if interest is compounded semi-annually?


Submit
Answer format: Percentage Round to: 4 decimal places (Example: 9.2434%, % sign required. Will accept decimal format rounded to 6 decimal places (ex: 0.092434))
Answer #1

Value of stock = Expected Dividend next year/(Required rate of return – Growth rate)

= 2.01/(14.47%-3.75%)

= $18.75

Amount required on 75th birthday = 2,760,505/(1.05) = $2,629,052.38

Amount required on 65th birthday = 2,629,052.38/(1.05)10

= $1,614,010.10

Future value of Annuity = Periodic Amount*[{(1+r)n – 1}/r]

Let annual deposits be x

1,614,010.10 = x*[{(1.05)40 -1}/0.05]

1,614,010.10 = 120.79977x

X = $13,361.04

Hence, annual contribution required = $13,361.04

Effective Annual rate = (1+periodic rate)^number of periods – 1

= (1+4%)^2 – 1

= 8.16%